OpenStudy (anonymous):

With these dimensions: Dimensions. The pool will hold 2,000 ft3 of water, be 8 inches deep(height), have a width of 5 feet and a length of 50 feet. What is the lateral area? Be in mind this is for a pool and the top side is missing.

OpenStudy (radar):

You will only need to compute the area of the 4 sides, as lateral area excludes the area of the base.

OpenStudy (radar):

Convert the 8 inch to dimension to feet: 8/12=2/3 ft. The width sides would be 2/3 ft. X 5 ft, Area A=2/3 X 5 since there are two sides the area of both sides: A= 2(2/3)(5) Compute the area of the length sides: A=(2/3)(50), again there are 2 long sides so the area for both long sides would be: A=(2)(2/3)(50) total area would be the sum of the 4 sides: A(total)=100(2/3) + 10(2/3): 200 20 220 ---- + ----- = ----- = 73 1/3 Sq. Ft. 3 3 3

OpenStudy (radar):

If decimal answer is desired then: 73.33333333333333333333333333 Sq ft.

OpenStudy (radar):

Just kidding, 73.33 ought to get it.

OpenStudy (anonymous):

Thanks! haha

OpenStudy (radar):

You're welcome and good luck with your studies.